A wire is 10m long and reseistance of wire is 20ohm. A potential difference of 10^-5volt/cm is to be created along the length of the wire . Find the value of R needed if Es=1.082V and the galvanometer shows no deflection.


1
Expert's answer
2021-05-31T16:44:02-0400

Gives

Length (l)=10 m

Resistance (R)=20 ohm

Potential difference (V)="10^{-3} V\/m"

Es=1.082V

R'=?

We know that V=IR

Current (I)="\frac{10^{-3}}{20}=5\times10^{-5}A"

Then we know that

"E_s=\frac{I(R'+R)}{l}"

Put value

"1.082=\frac{5\times10^{-5}\times(R'+20)}{10}"

R'=216380 ohm

R'=0.22"M\Omega"